# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ IV.4 Impact of the radiative energy loss on nPDF evaluation
As discussed in Section III.1, nPDFs are extracted from a variety of QCD processes in lepton-nucleus and proton-nucleus collisions. Over the last few years, it has been shown that the production of open-charm mesons and charmonia in $p+A$ collisions at LHC energies provides significant constraints on extraction of nPDFs, especially in the still poorly constrained gluon sector at small $x$ [223, 4, 224, 8]. These processes, however, are also sensitive to medium-induced gluon radiation and, hence, to fully coherent energy loss [13, 381, 231]. Thus, some fraction of the $A$ dependence of $D$ meson and $J/\psi$ production can be attributed to radiative energy loss and should therefore not be absorbed into the nPDFs. Despite the good description of world data in current global fits, analyses based on QCD factorization and neglecting factorization-breaking effects such as FCEL may lead to a biased nPDF extraction.
@@
This challenges the current strategy for extracting nPDFs from global fits. To circumvent this issue, one possibility is to restrict the fits to observables insensitive to FCEL, such as electroweak bosons, Drell-Yan, and, to a lesser extent, jet production in $p+A$ collisions, in addition to DIS measurements. This, however, weakens the nPDF constraints, especially for small-$x$ gluons, as illustrated by the purple band in Fig. 16, showing the nNNPDF3.0_noLHCbD set, which excludes the LHCb $D$ meson data.
@@
An alternative strategy would be to include the heavy flavor data in global nPDF fits, but only after incorporating FCEL effects into the analyses. This approach was recently explored in a case study [409], where $J/\psi$ production was computed at leading order in the color evaporation model with several nPDF sets included in the $p+A$ calculations: EPPS21 [4], nCTEQ15WZSIH [224], and nNNPDF3.0_noLHCbD [8]. Preliminary results are shown in Fig. 16 for a Bayesian reweighting of the nNNPDF3.0_noLHCbD set, using ALICE $J/\psi$ measurements [97] and assuming either the presence (blue band) or absence (orange band) of FCEL. In both cases, the $J/\psi$ data strongly constrain the gluon nPDF at small $x$, confirming earlier findings [223, 224]. The figure shows that the gluon nPDF ratio using the $J/\psi$ data with FCEL included is clearly separated from the ratio obtained with the $J/\psi$ data and no FCEL. Without FCEL, the gluon nPDF ratio is approximately $R_{g}\gtrsim 0.6$ at $x\sim 10^{-5}$, similar to the nNNPDF3.0 set which includes LHCb $D$ meson data (not shown). In contrast, $R_{g}\simeq 0.9$ when FCEL is taken into account. [^8] In this example, the transport coefficient at $x=10^{-2}$ is taken to be $\hat{q}_{0}=0.07$ GeV2/fm. A more systematic and in-depth investigation is in progress.
@@
In conclusion, nPDF sets reweighted with heavy flavor data are shown to be very sensitive to the inclusion of FCEL effects, demonstrating that taking FCEL into account can significantly impact the extraction of the gluon nPDF in particular. This highlights the need to consistently implement such effects in global fits incorporating heavy-flavor probes of $h+A$ collisions in order to determine the correct level of small-$x$ leading-twist gluon shadowing.
@@ IV.5 Energy loss at small $x$
The medium transport parameters (interaction range $\nu$, mean free path $\lambda$, or transport coefficient $\hat{q}$) must be sufficiently large for CNM energy loss to be significant. In a microscopic picture, this, in turn, implies that the nuclear target must be in a dense partonic system. In the saturation picture, such a dense state can be probed either by probing large $A$ at intermediate $x$ (the MV regime) or at small $x$ through BK-JIMWLK evolution. The relevant quantity is the coherence length (longitudinal wavelength of the radiated gluon), inversely proportional to the $x$. If the coherence length is larger than the nuclear size, this FCEL is contained in NLO corrections to the CGC [410, 303]. On the other hand, if the coherence length of the radiation is larger than the nucleon size but smaller than the nuclear size, we have “partially coherent” energy loss. This could be a serious issue for the CGC because the treatment of this regime from first principles is unknown in the saturation framework. Much of our knowledge of gluon saturation in nuclei comes from forward rapidity RHIC data for $p\,({\rm d})+{\rm Au}$ collisions where, due to the steeply falling $p_{T}$ spectra, even a small contribution from energy loss could result in a major modification of the spectra. While this is not relevant for total cross sections (structure functions) it could play a significant role for forward rapidity observables. How to separate energy loss effects from suppression generated by genuine gluon saturation dynamics in regions of phase space where both can contribute is an open question.
@@ IV.6 Observables
  • Colorless processes either in the initial-state or final-state could be particularly relevant because they are insensitive to the fully coherent regime and thus solely be sensitive to LPM energy loss and nPDF effects. This is the case for hadron production in low energy semi-inclusive DIS (SIDIS) on nuclei [393] where propagating quarks lose energy in the final state before fragmenting into hadrons. In addition, normalizing the hadrons yields in SIDIS by the number of DIS events drastically reduce the impact of nPDF effects. The Drell-Yan process in $\pi+A$ and $p+A$ collisions is also particularly relevant. At high energies, LPM energy loss effects on Drell-Yan should be negligible, making it a clean nPDF probe [411]. Conversely, nPDF effects in low energy $\pi+A$ and $p+A$ collisions are either small or tightly constrained by DIS measurements, making the Drell-Yan process well suited for studies of LPM energy loss in the large $x_{F}$ regime [378, 41, 379, 42, 401].
  • FCEL is not universal effect because it depends on the color structure of the partonic process. Thus FCEL cannot be absorbed into the wavefunction of the target nucleus. This non-universality could best be seen by comparing measurements in $e+A$ and $p+A$ collisions: the $A$ dependence of hadron production should be affected by FCEL only in $p+A$. However, nPDF effects should be similar in both systems, provided the typical momentum fraction probed in the target nucleon, $x\simeq(M^{2}+Q^{2})/(W^{2}+Q^{2})$ and $x_{2}\simeq M/\sqrt{s}\,e^{-y}$ respectively for the production of a particle with mass $M$, are equal. Specifically, one would expect the nuclear production ratio of, e.g. $J/\psi$ production to be [^9] Assuming LPM energy loss to be small for high energy $J/\psi$ production.
@@
    $R_{pA}^{J/\psi}(y,x_{2})$ $\simeq$ $R_{pA}^{\text{nPDF+FCEL}}(y,x_{2})\approx R_{g}^{\text{A}}(x_{2})$ (10)
    ×$\times R_{pA}^{\text{FCEL}}(y,x_{2})\,.$
@@
  • There is a strong correlation between the small-$x$ gluon and sea quark distributions. Reference [411] showed that the double ratio $R_{pA}^{J/\psi}/R_{pA}^{\text{DY}}$ at forward rapidity was close to or somewhat above unity for the available nPDF sets [411] but significantly smaller than unity for FCEL. This observable could therefore help separate FCEL from nPDF effects in this region.
@@
  • The observed $p_{T}$ broadening, $\Delta p_{T}^{2}$ in Drell-Yan and quarkonium production or SIDIS can be used to extract the transport coefficient $\hat{q}$ more precisely and thus and provide constraints on energy loss. This observable is mainly independent of nPDF and energy loss effects at low transverse momentum. Indeed, in this kinematic, the dependence of the transverse momentum cross section on the nuclear PDFs and energy loss is small; $R_{hA}^{\text{nPDF+E.loss}}(p_{T}\lesssim M)$ remains approximately constant. As a result, these effects tend to cancel out in the calculation of $\Delta p_{T}^{2}$ (see Ref. [15] for further discussion).
@@ V.1 Collisional dissociation of open heavy flavor and quarkonia
Collisional and radiative parton-level processes can be complemented by hadron-level interactions. Heuristically, the physics picture that underlies such models is easy to understand. The time to form a hadron in its rest frame is on the order of its size, $\sim 1$ fm which must be boosted to the lab frame by the Lorentz $\gamma$ factor. At low energies, even light hadrons can have short formation times. This fact was exploited in a model of hadron absorption [413, 414] applied to HERMES fixed target $e+A$ measurements. For example, if a finite mass parton, $m_{Q}$, fragments into a hadron of mass $m_{h}$ and a light parton, the formation length, $\Delta y^{+}\sim\tau_{\rm form}$, where $\tau_{\rm form}$ is the formation time, is, [415],
@@
The hadron and parton mass dependence in Eq. (11) suggests that, for open heavy flavor [415, 418] and quarkonia [419, 420], the formation length can be considerably shorter than the nuclear size, $\Delta y^{+}<2R_{A}$, even for large boosts. This further implies that in-medium hadron-level dynamics may be dominant at small and moderate $p_{T}$. A microscopic model that evaluates the dissociation rate from the distortion of the meson wave function due to collisional interactions in matter has been formulated and phenomenologically applied to successfully describe the suppression patterns of: $D$ and $B$ mesons; their semileptonic decays; and and quarkonium suppression in $A+A$ collisions.
@@
Figure 17: Top: Simplified calculations of parton energy loss (solid lines) and absorption within nuclei (dashed lines) on the momentum fraction $z_{h}$ distribution of hadrons in the acceptance of the forward silicon tracker at the EIC. The projected statistical precision (points) of forward rapidity $B$ and $D$ meson measurements with integrated luminosities in $e+p$ ($e+{\rm Au}$) collisions is 10 fb-1 (500 nb-1) illustrates the power of the EIC to discriminate between energy loss and hadron absorption [398]. Bottom: Comparison between the suppression of several charmonium states in CNM (dashed lines) and the QGP (solid lines) due to dissociation. Adapted from Ref. [421].
@@
Similar formation considerations apply to cold nuclear matter and deep inelastic scattering on nuclei. On general grounds, one would expect that the larger center-of-mass energies at the EIC compared to HERMES will lead to larger boosts so that the nuclear modification factor for light hadrons, such as pions and kaons, will be predominantly driven by energy loss. Nevertheless, this remains an open question that cannot currently be unambiguously answered. The characteristic shape of the charm and bottom quark fragmentation functions may be key to identifying the origin of CNM meson suppression [398]. As shown on the left-hand side of Fig. 17, absorption always leads to a reduction of the differential cross section as a function of $z$. In contrast, in-medium parton shower modifications or energy loss lead to enhanced of $D$ and $B$production at intermediate and small $z$. In addition, the narrower, bell-shaped, $B$ meson fragmentation function, shifted to larger $z$, signifying that less of the $b$ quark energy is lost to fragmentation, is responsible for their larger modification relative to $D$ mesons. It is expected, as shown by the simulated error bars for projected yearly luminosities, that this physics can be studied at the EIC.
@@
In contrast to open heavy flavor, at the parton level, quarkonia originate from a heavy $Q\overline{Q}$ pair produced in a specific color and angular momentum state at very early times that are inversely proportional to the virtuality of the hard process. Consequently, interactions of the heavy quarks with the medium start early and the energy dependence of the suppression is very weak [421]. A preliminary comparison of the suppression of $J/\psi$, $\psi$(2S), and $\chi_{c}$ in CNM, $R_{eA}$ and the QGP, $R_{AA}$, is shown on the right-hand side of Fig. 17. While weakly bound states like the $\psi$(2S) and $\chi_{c}$ are more suppression in the QGP due to thermal wavefunction effects, the CNM and QGP effects on the $J/\psi$ are similar. This is because the thermal screening of the potential for the most tightly bound $Q\overline{Q}$ states is very small, while collisional dissociation depends on the interplay between the quarkonium binding and the space-time dynamics of the nuclear matter transport coefficients. In all cases, the nuclear modification is significant, a factor of two or more, and should be easily observable experimentally. The future EIC will provide the first opportunity to study quarkonium modification in $e+A$ collisions.
@@ V.2.1 Absorption by Nucleons
If it is produced as a small, growing color singlet until reaching its formation time when it becomes a fully formed bound state. In this case, absorption is assumed to only play a role if the state can be formed inside the target. Thus, at most energies, the state is only produced inside the target, where absorption is effective, at negative $x_{F}$. The larger the state, the longer its formation time and the larger its asymptotic absorption cross section, assuming $\sigma_{\rm abs}$ grows as the square of the quarkonium radius.
If, on the other hand, the proto-charmonium state, $|(c\overline{c})_{8}g\rangle$, is assuming to be in color single through its characteristic octet lifetime, all the charmonium states pass through matter with the constant absorption cross section, particularly at forward $x_{F}$. At negative $x_{F}$, the octet path length is shorter and it can neutralize its color and be absorbed as a color singlet with a different cross section. Here, the $(c\overline{c}))_{8}$ configuration is not yet on shell.
Finally, in nonrelativistic QCD, the charmonium states are an admixture of singlet and octet production. In this case, the color singlet dominated $\chi_{c}$ states would have a near linear $A$ dependence at forward $x_{F}$ while the $J/\psi$ and $\psi^{\prime}$, dominated by octet production, would have a less than linear $A$ dependence at forward $x_{F}$. In this case, the quarkonium states would come on shell after the characteristic formation time for color singlet states and after the octet lifetime for color octet states. Before these characteristic times, the $c\overline{c}$ pair interacts with the nuclear matter.
All of these scenarios were examined in detail in Ref. [424]. Unfortunately the $\chi_{c}$ $A$ dependence has not yet been measured with precision, a measurement that could distinguish between types of production and absorption.
@@
The extraction of $\sigma_{\text{abs}}$ is affected by both nPDF and FCEL contributions. When $R_{hA}^{\text{nPDF}}\lesssim 1$ (in the EMC or shadowing) regions, $\sigma_{\text{abs}}$ may be underestimated, if $R_{hA}^{\text{nPDF}}\gtrsim 1$ (in the antishadowing region), absorption could be overestimated. Moreover, if FCEL is not properly distinguished from absorption, $\sigma_{\text{abs}}$ may also be overestimated because $R_{hA}^{\text{FCEL}}<1$. The challenge here is to establish observables that either nPDF or FCEL effects, as discussed in Sec. II and IV, or provide a comprehensive description of the charmonium production cross section.

### v2

@@ IV.4 Impact of the radiative energy loss on nPDF evaluation
As discussed in Section III.1, nPDFs are extracted from a variety of QCD processes in lepton-nucleus and proton-nucleus collisions. Over the last few years, it has been shown that the production of open-charm mesons and charmonia in $p+A$ collisions at LHC energies provides significant constraints on nPDF extractions, especially in the still poorly constrained gluon sector at small $x$ [225, 4, 226, 8]. These processes, however, are also sensitive to medium-induced gluon radiation and, hence, to fully coherent energy loss [13, 383, 233]. Thus, some fraction of the $A$ dependence of $D$ meson and $J/\psi$ production can be attributed to radiative energy loss and should therefore not be absorbed into the nPDFs. Despite the good description of world data in current global fits, analyses based on QCD factorization and neglecting factorization-breaking effects such as FCEL may lead to a biased nPDF extraction.
@@
This challenges the current strategy for extracting nPDFs from global fits. To circumvent this issue, one possibility is to restrict the fits to observables insensitive to FCEL, such as electroweak bosons, Drell-Yan, and, to a lesser extent, jet production in $p+A$ collisions, in addition to DIS measurements. This, however, weakens the nPDF constraints, especially for small $x$ gluons, as illustrated by the purple band in Fig. 16, showing the nNNPDF3.0_noLHCbD set, which excludes the LHCb $D$ meson data.
@@
An alternative strategy would be to include the heavy flavor data in global nPDF fits, but only after incorporating FCEL effects into the analyses. This approach was recently explored in a case study [411] where $J/\psi$ production was computed at leading order in the color evaporation model with several nPDF sets included in the $p+A$ calculations: EPPS21 [4], nCTEQ15WZSIH [226], and nNNPDF3.0_noLHCbD [8]. Preliminary results are shown in Fig. 16 for a Bayesian reweighting of the nNNPDF3.0_noLHCbD set, using ALICE $J/\psi$ measurements [97] and assuming either the presence (blue band) or absence (orange band) of FCEL. In both cases, the $J/\psi$ data strongly constrain the gluon nPDF at small $x$, confirming earlier findings [225, 226]. The figure shows that the gluon nPDF ratio using the $J/\psi$ data with FCEL included is clearly separated from the ratio obtained with the $J/\psi$ data and no FCEL. Without FCEL, the gluon nPDF ratio is approximately $R_{g}\gtrsim 0.6$ at $x\sim 10^{-5}$, similar to the nNNPDF3.0 set which includes LHCb $D$ meson data (not shown). In contrast, $R_{g}\simeq 0.9$ when FCEL is taken into account. [^8] In this example, the transport coefficient at $x=10^{-2}$ is taken to be $\hat{q}_{0}=0.07$ GeV2/fm. A more systematic and in-depth investigation is in progress.
@@
In conclusion, nPDF sets reweighted with heavy flavor data are shown to be very sensitive to the inclusion of FCEL effects, demonstrating that taking FCEL into account can significantly impact the extraction of the gluon nPDF in particular. This highlights the need to consistently implement such effects in global fits incorporating heavy-flavor probes of $h+A$ collisions in order to determine the correct level of small $x$ leading-twist gluon shadowing.
@@ IV.5 Energy loss at small $x$
The medium transport parameters (interaction range $\mu$, mean free path $\lambda$, or transport coefficient $\hat{q}$) must be sufficiently large for CNM energy loss to be significant. In a microscopic picture, this, in turn, implies that the nuclear target must be a dense system of partons. In the saturation picture, such a dense state can be probed either by probing large $A$ at intermediate $x$ (the MV regime) or at small $x$ through BK-JIMWLK evolution. The relevant quantity is the coherence length (longitudinal wavelength of the radiated gluon), inversely proportional to $x$. If the coherence length is longer than the nuclear size, this FCEL is contained in NLO corrections to the CGC [412, 305]. On the other hand, if the coherence length of the radiation is larger than the nucleon size but smaller than the nuclear size, we have “partially coherent” energy loss. This could be a serious issue for the CGC because the treatment of this regime from first principles is unknown in the saturation framework. Much of our knowledge of gluon saturation in nuclei comes from forward rapidity RHIC data for $p\,({\rm d})+{\rm Au}$ collisions where, due to the steeply falling $p_{T}$ spectra, even a small energy loss contribution could result in major modification of the spectra. While this is not relevant for total cross sections (structure functions) it could play a significant role for forward rapidity observables. How to separate energy loss effects from suppression generated by genuine gluon saturation dynamics in regions of phase space where both can contribute is an open question.
@@ IV.6 Observables
  • Colorless processes either in the initial-state or final-state could be particularly relevant because they are insensitive to the fully coherent regime and thus solely be sensitive to LPM energy loss and nPDF effects. This is the case for hadron production in low energy semi-inclusive DIS (SIDIS) on nuclei [395] where propagating quarks lose energy in the final state before fragmenting into hadrons. In addition, normalizing the hadron yields in SIDIS by the number of DIS events drastically reduce the impact of nPDF effects. The Drell-Yan process in $\pi+A$ and $p+A$ collisions is also particularly relevant. At high energies, LPM energy loss effects on Drell-Yan should be negligible, making it a clean nPDF probe [413]. Conversely, nPDF effects in low energy $\pi+A$ and $p+A$ collisions are either small or tightly constrained by DIS measurements, making the Drell-Yan process well suited for studies of LPM energy loss in the large $x_{F}$ regime [380, 41, 381, 42, 403].
  • FCEL is not universal effect because it depends on the color structure of the partonic process. Thus FCEL cannot be absorbed into the wavefunction of the target nucleus. This non-universality could best be seen by comparing measurements in $e+A$ and $p+A$ collisions: the $A$ dependence of hadron production should be affected by FCEL only in $p+A$. However, nPDF effects should be similar in both systems, provided the typical momentum fractions probed in the target nucleon are equal: $x\simeq(M^{2}+Q^{2})/(W^{2}+Q^{2})$ in $e+A$ and $x_{2}\simeq M/\sqrt{s}\,e^{-y}$ in $p+A$ for production of a particle with mass $M$. Specifically, one would expect the nuclear production ratio of, e.g. $J/\psi$ production to be [^9] Assuming LPM energy loss to be small for high energy $J/\psi$ production.
@@
    $R_{pA}^{J/\psi}(y,x_{2})$ $\simeq$ $R_{pA}^{\text{nPDF+FCEL}}(y,x_{2})$ (9)
    $\approx$ $R_{g}^{\text{A}}(x_{2})\times R_{pA}^{\text{FCEL}}(y,x_{2})\,.$ (10)
@@
  • There is a strong correlation between the small $x$ gluon and sea quark distributions. Reference [413] showed that the double ratio $R_{pA}^{J/\psi}/R_{pA}^{\text{DY}}$ at forward rapidity was close to or somewhat above unity for the available nPDF sets [413] but significantly smaller than unity for FCEL. This observable could therefore help separate FCEL from nPDF effects in this region.
@@
  • The observed $p_{T}$ broadening, $\Delta p_{T}^{2}$ in Drell-Yan and quarkonium production or SIDIS can be used to extract the transport coefficient $\hat{q}$ more precisely and thus provide constraints on energy loss. This observable is mainly independent of nPDF and energy loss effects at low $p_{T}$. Indeed, in this kinematic regime, the dependence of $d\sigma/dp_{T}$ on nPDFs and energy loss is small; $R_{hA}^{\text{nPDF+E.loss}}(p_{T}\lesssim M)$ remains approximately constant. As a result, these effects tend to cancel out in the calculation of $\Delta p_{T}^{2}$ (see Ref. [15] for further discussion).
@@ V.1 Collisional dissociation of open heavy flavor and quarkonia
Collisional and radiative parton-level processes can be complemented by hadron-level interactions. Heuristically, the physics picture that underlies such models is easy to understand. The time to form a hadron in its rest frame, $\tau_{\rm form}$ is on the order of its size, $\sim 1$ fm, which must be boosted to the lab frame by the Lorentz $\gamma$ factor. At low energies, even light hadrons can have short formation times. This fact was exploited in a model of hadron absorption [415, 416] applied to HERMES fixed target $e+A$ measurements. For example, if a finite mass parton, $M_{Q}$, fragments into a hadron of mass $M_{h}$ and a light parton, the formation length, $\Delta y^{+}\sim\tau_{\rm form}$, is [417]
@@
The hadron and parton mass dependence in Eq. (11) suggests that, for open heavy flavor [417, 420] and quarkonia [421, 422], the formation length can be considerably shorter than the nuclear size, $\Delta y^{+}<2R_{A}$, even for large boosts. This further implies that in-medium hadron-level dynamics may be dominant at small and moderate $p_{T}$. A microscopic model that evaluates the dissociation rate from the distortion of the meson wave function due to collisional interactions in matter has been formulated and phenomenologically applied to successfully describe the suppression patterns of $D$ and $B$ mesons; their semileptonic decays; and quarkonium suppression in $A+A$ collisions.
@@
Figure 17: Top: Simplified calculations of parton energy loss (solid lines) and absorption within nuclei (dashed lines) on the momentum fraction $z_{h}$ distribution of hadrons in the acceptance of the forward silicon tracker at the EIC. The projected statistical precision (points) of forward rapidity $B$ and $D$ measurements with integrated luminosities in $e+p$ ($e+{\rm Au}$) collisions is 10 fb-1 (500 nb-1) illustrates the power of the EIC to discriminate between energy loss and absorption [400]. Bottom: Comparison between the suppression of several charmonium states in CNM (dashed lines) and the QGP (solid lines) due to dissociation. Adapted from Ref. [423].
@@
Similar formation considerations apply to CNM and nuclear DIS. On general grounds, one would expect that the larger center-of-mass energies at the EIC relative to HERMES will lead to larger boosts so that the nuclear modification factor for light hadrons, such as pions and kaons, will be predominantly driven by energy loss. Nevertheless, this remains an open question that cannot currently be unambiguously answered. The characteristic shape of the charm and bottom quark fragmentation functions may be key to identifying the origin of CNM meson suppression [400]. As shown on the top panel of Fig. 17, absorption always leads to a reduction of the differential cross section as a function of $z$. In contrast, in-medium parton shower modifications or energy loss lead to enhanced $D$ and $B$ production at intermediate and small $z$. In addition, the narrower $B$ meson fragmentation function is shifted to larger $z$, signifying that less $b$ quark energy is lost to fragmentation. This shift is responsible for the larger $B$ modification relative to $D$ mesons. It is expected, as indicated by the simulated error bars for projected yearly luminosities, that this physics can be studied at the EIC.
@@
In contrast to open heavy flavor, at the parton level, quarkonia originate from a heavy $Q\overline{Q}$ pair produced in a specific color and angular momentum state at very early times, inversely proportional to the virtuality of the hard process. Consequently, interactions of the heavy quarks with the medium start early and the energy dependence of the suppression is very weak [423]. A preliminary comparison of the suppression of $J/\psi$, $\psi$(2S), and $\chi_{c}$ in CNM, $R_{eA}$, and in the QGP, $R_{AA}$, is shown on the bottom panel of Fig. 17. While weakly bound states like the $\psi$(2S) and $\chi_{c}$ are more suppressed in the QGP due to thermal wavefunction effects, the CNM and QGP effects on the $J/\psi$ are similar. This is because the thermal screening of the potential for the most tightly bound $Q\overline{Q}$ states is very small, while collisional dissociation depends on the interplay between the quarkonium binding and the space-time dynamics of the nuclear matter transport coefficients. In all cases, the nuclear modification is significant, a factor of two or more, and should be easily observable experimentally. The future EIC will provide the first opportunity to study quarkonium modification in $e+A$ collisions.
@@ V.2.1 Absorption by Nucleons
If it is produced as a small color singlet, it grows until reaching its formation time when it becomes a fully formed bound state. In this case, absorption is assumed to only play a role if the state can be formed inside the target. Thus, at most energies, the state is only produced inside the target, where absorption is effective, at negative $x_{F}$. The larger the state, the longer its formation time and the larger its asymptotic absorption cross section, assuming $\sigma_{\rm abs}$ grows as the square of the quarkonium radius.
If, on the other hand, the proto-charmonium state, $|(c\overline{c})_{8}g\rangle$, is assumed to be in color singlet through its characteristic octet lifetime, all the charmonium states pass through matter with a constant absorption cross section, particularly at forward $x_{F}$. At negative $x_{F}$, the octet path length is shorter and it can neutralize its color and be absorbed as a color singlet with a different cross section. Here, the $(c\overline{c})_{8}$ configuration is not yet on shell.
Finally, in nonrelativistic QCD, the charmonium states are an admixture of singlet and octet production. In this case, the color singlet dominated $\chi_{c}$ states would have a near linear $A$ dependence at forward $x_{F}$ while the $J/\psi$ and $\psi$(2S), dominated by octet production, would have a less than linear $A$ dependence at forward $x_{F}$. In this case, the quarkonium states would come on shell after the characteristic formation time for color singlet states and after the octet lifetime for color octet states. Before these characteristic times, the $c\overline{c}$ pair interacts with the nuclear matter.
All of these scenarios were examined in detail in Ref. [426]. Unfortunately the $\chi_{c}$ $A$ dependence has not yet been measured with precision, a measurement that could distinguish between types of production and thus also of absorption.
@@
The extraction of $\sigma_{\text{abs}}$ is affected by both nPDF and FCEL contributions. When $R_{hA}^{\text{nPDF}}\lesssim 1$ (in the EMC or shadowing) regions, $\sigma_{\text{abs}}$ may be underestimated, if $R_{hA}^{\text{nPDF}}\gtrsim 1$ (in the antishadowing region), absorption could be overestimated. Moreover, if FCEL is not properly distinguished from absorption, $\sigma_{\text{abs}}$ may also be overestimated because $R_{hA}^{\text{FCEL}}<1$.